The sum of ages of Caroline and Izzy is 19 years old. In three years, 6 times Izzy's age will equal 4 times Caroline's age. How old is each now?
Answer by Boreal(15235)

x=Caroline
19-x=Izzy.
[(19-x)+3]*6=4(x+3)
(22-x)*6=4x+12
132-6x=4x+12
120=10x
x=12 Caroline is 12
Izzy is 7
In 3 years, Caroline will be 15 and Izzy 10. Six times Izzy's age is 60-4 times Caroline's age.